### v1

# Onsager vortex clusters on a sphere

Abstract
We study Onsager vortex clustered states in a shell-shaped superfluid containing a large number of quantum vortices. In the incompressible limit and at low temperatures, the relevant problem can be boiled down to the statistical mechanics of neutral point vortices confined on a sphere. We analyze rotation free vortex clustered states within the mean field theory in the microcanonical ensemble. We find that the sandwich state, which involves the separating of vortices with opposite circulation and the clustering of vortices with same circulation around the poles and the equator, is the maximum entropy vortex distribution, subject to zero angular momentum constraint. The dipole momentum vanishes for the sandwich state and the quadrupole tensor serves as an order parameter to characterize the vortex cluster structure. For given finite angular momentum, the equilibrium vortex distribution forms a dipole structure, i.e., vortices with opposite sign are separated and are accumulated around the south and north pole, respectively. The conditions for the onset of clustering, and the exponents associated with the quadrupole moment and the dipole moment as functions of energy, are obtained within the mean field theory. At large energies, we obtain asymptotically exact vortex density distributions using the stereographic projection method, which give rise the parameter bounds for the vortex clustered states. The analytical predictions are in excellent agreement with microcanonical Monte Carlo simulations.
I Introduction
Coherent large vortex structures can occur in various bounded flows, examples range from Great Red Spot in Jupiter’s atmosphere Vasavada et al. (1998), to giant vortex clusters in atomic Bose-Einstein condensates (BECs) Gauthier et al. (2019); Johnstone et al. (2019), and in dissipative quantum fluids of exciton-polaritons Panico et al. (2023). For a vortex system in a bounded domain, the phase space per vortex is also bounded. The formation of vortex clusters is due to that the system favors the clustering of like-sign vortices at high energies Onsager (1949); Eyink and Sreenivasan (2006). Onsager formulates the statistical mechanics of a bounded point vortex system in the microscopic ensemble and demonstrates that the clustered state is at negative absolute temperature. Since then clustering of vortices on bounded domains has been investigated extensively Joyce and Montgomery (1973); Edwards and Taylor (1974); Williamson (1977); Pointin and Lundgren (1976); Campbell and O’Neil (1991); Smith (1989); Smith and O’Neil (1990); Kiessling (1995); Yatsuyanagi et al. (2005); Esler and Ashbee (2015); Qi and Marston (2014); Siggia and Aref (1981); Reeves et al. (2012); Billam et al. (2015); Simula et al. (2014); Gurarie and Chow (2004); Kiessling and Wang (2012); Yu et al. (2016); Salman and Maestrini (2016); Valani et al. (2018); Han and Tsubota (2018, 2019); Skaugen and Angheluta (2016, 2017); Kanai and Guo (2021); Patil and Dunkel (2021); Pandit et al. (2017); Chavanis and Sommeria (1996); Chavanis (2002); Xiong et al. (2023). The clustering of vortices is also closely related to the end state of inverse energy cascades Billam et al. (2014); Skaugen and Angheluta (2016); Reeves et al. (2013), which involves energy transport from small to large scales in two-dimensional (2D) turbulence.
Previous studies on Onsager vortices mainly focus on systems confined on a flat region. Recent experimental realizations of BECs Aveline et al. (2020) and ultracold atomic bubbles Carollo et al. (2022) in International Space Station open a possibility to investigate a bubble trapped superfluid and vortex physics on curved surfaces experimentally. Curvature and topology of a surface have non-trivial effects on BECs Tononi and Salasnich (2019); Tononi et al. (2020); Tononi and Salasnich (2023), vortex dynamics Bogomolov (1978); CROWDY (2004); Hally (1980); Polvani and Dritschel (1993); Kimura (1999); Newton (2001); Alobaidi et al. (2006); Aref (2007); Vitelli and Turner (2004); Turner et al. (2010); Dritschel and Boatto (2015); Dritschel et al. (2015); Doshi and Gromov (2021); Bereta et al. (2021); Caracanhas et al. (2022); Padavić et al. (2020); Saito and Hayashi (2023); Li and Efimkin (2023); Xiong and Yu (2024), and the Berezinskii-Kosterlitz-Thouless (BKT) transition Tononi et al. (2022). The formation and structure of Onsager vortices also strongly depend on the surface geometry. Numerical simulations of the Gross-Pitaevskii equation show that, different from the giant dipole configuration for vortices confined to a discs Simula et al. (2014); Yu et al. (2016), in a spherical shell trap uniformly distributed neutral vortices form a quadruple cluster through the evaporative heating process Kanai and Guo (2021). However, it is not clear whether the quadruple clustered state is the statistically most favored state for given energy, angular momentum and vortex number.
In this paper, we systematically investigate rotation-free clustered states of a neutral vortex system confined on a sphere. We extend the mean field theory of vortex statistical mechanics developed by Joyce and Montgomery Joyce and Montgomery (1973) to a closed surface. In particular, we analyze the self-consistent equation of the stream function on a sphere near the uniform state and in the high energy limit. We find that, under constraint of zero angular momentum, the sandwich state where vortices with opposite sign are distributed around the poles and the equator respectively, is the maximum entropy state and has higher statistical weight than the quarupole state. This theoretical prediction is conformed by Monte Carlo simulations.
II Point vortices on a closed surface
For a superfluid on a plane, the circulation of a quanutm vortex is quantized in units of circulation quantum $\kappa\equiv 2\pi\hbar/m$ Pitaevskii and Stringari (2016), and the vorticity has a singularity at the vortex core $\mathbf{r}_{i}$: $\omega(\mathbf{r})=\nabla\times\mathbf{u}=\kappa_{i}\delta(\mathbf{r}-\mathbf{r}_{i})=\kappa\sigma_{i}\delta(\mathbf{r}-\mathbf{r}_{i})$ with sign $\sigma_{i}=\pm 1$ for singly charged vortices. Here $m$ is the atomic mass and $\mathbf{u}$ is the superfluid velocity. The quantization ensures that the vorticity of a quantum vortex concentrates around the core region Eyink and Sreenivasan (2006). Hence when the mean separation between quantum vortices $\ell$ is much larger than the vortex core size $\xi$ Fetter (1998); Barenghi and Parker (2016), the point vortex model (PVM) describes well the dynamics of quantum vortices in the incompressible limit Onsager (1949); Eyink and Sreenivasan (2006); Novikov (1975); Aref et al. (1999), provided vortex annihilation can be neglected. The PVM also describes 2D superfluid transition at finite temperature, which is the celebrated BKT transition Kosterlitz and Thouless (1973); Kosterlitz (1974).
On curved surfaces, a covariant PVM has been well formulated Hally (1980); Kimura (1999); Dritschel and Boatto (2015). Here we summarize the main results of the PVM on a closed surface. In isothermal coordinates, the metric on the surface is
$ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}=h(x^{1},x^{2})[(dx^{1})^{2}+(dx^{2})^{2}],$ (1)
where $g_{12}=g_{21}=0$, $g_{11}=g_{22}=h(x^{1},x^{2})$, and $h(x^{1},x^{2})$ is a positive function and exists locally for 2D surfaces Chern (1955). For incompressible fluids, the component fluid velocity $u_{a}=\epsilon_{a}^{\,\,\,b}\nabla_{b}\psi$, where $\psi$ is the stream function, $\nabla_{b}$ is the Levi-Civita covariant derivative, $\epsilon^{\beta}_{\,\,\,\,\alpha}=g^{\beta\gamma}\epsilon_{\gamma\alpha}$, $\epsilon_{\gamma\alpha}=\sqrt{\det g_{\mu\nu}}\,\tilde{\epsilon}_{\gamma\alpha}$ is the Levi-Civita tensor and $\tilde{\epsilon}_{\gamma\alpha}$ is the Levi-Civita symbol. The vorticity $\omega=\epsilon^{\mu}_{\,\,\,\,\nu}\nabla_{\mu}u^{\nu}=-\Delta\psi$, where $\Delta\equiv\nabla^{\mu}\nabla_{\mu}=(1/\sqrt{g})\partial_{i}\left(\sqrt{g}g^{ij}\partial_{j}\right)=h^{-1}(x^{1},x^{2})(\partial^{2}_{x^{1}}+\partial^{2}_{x^{2}})$ is the Laplace-Bertrami operator and $g=\det g_{ij}$.
Let us also introduce complex coordinates $z=x^{1}+ix^{2}$, $\partial_{z}\equiv(\partial_{1}-i\partial_{2})/2$, $\partial_{\bar{z}}\equiv(\partial_{1}+i\partial_{2})/2$ and complex velocity $u\equiv u^{1}-iu^{2}$. In complex notations, $u=2ih^{-1}\partial_{z}\psi$ and the stream function
$\psi(z)=\kappa\sum_{i}\sigma_{i}G(z,z_{i}),$ (2)
where $z_{i}$ is the position of the $i$th vortex, $G(z,z_{i})$ is the Green’s function satisfying Dritschel and Boatto (2015)
$\Delta G(z,z_{i})=-\delta_{z,z_{i}}+1/\Omega,$ (3)
$\Omega$ is the area of the surface, and $\delta_{z,z_{k}}\equiv h^{-1}\delta(z-z_{k})$.
The Hamiltonian of point vortices reads
$H=-\frac{\rho_{s}}{2}\sum_{i\neq j}\kappa_{i}\kappa_{j}G(z_{i},z_{j})-\frac{\rho_{s}}{2}\sum_{i}\kappa^{2}_{i}R_{\rm robin}(z_{i}),$ (4)
where $\rho_{s}$ is the superfluid (mass) density,
$R_{\rm robin}(z)=\lim_{w\rightarrow z}\left[G(z,w)+\frac{1}{2\pi}\log d(z,w)\right]$ (5)
is the celebrated Robin function and $d(z,w)$ is the geodesic distance between points $w$ and $z$. When $w\rightarrow z$, $d(z,w)=ds+o(|w-z|)=h|w-z|+o(|w-z|)$ and only the leading order term contributes the Robin function.
The kinetic energy of the fluid is
[…]
where $H_{\text{singular}}=(\rho_{s}/4\pi)\sum_{i}\kappa^{2}_{i}\log d(z_{i},z_{i})$. For the point vortex model, $H_{\text{singular}}=\infty$. In practice, this term is regularized by the vortex core size $\xi$ which serves the ultraviolet cut-off of the PVM, and $H_{\text{singular}}=(\rho_{s}/4\pi)\sum_{i}\kappa^{2}_{i}\log\xi$. Hereafter, we set $\rho_{s}=\kappa=1$ for convenience.
III Statistical mechanics of point vortices on a closed surface
On a closed surface, the total vorticity must vanish, namely, $\sum_{i}\kappa_{i}=0$. We consider a system consisting of $N_{+}$ positive vortices and $N_{-}=N_{+}$ negative vortices. The total vortex number is $N=N_{+}+N_{-}$. The model Eq. (4) exhibits Onsager clustered states at high energies in a bounded domain. In clustered phases, the energy $H\sim N^{2}$, hence in order to describe clustered phases in the proper thermodynamics limit, we need to make the rescaling $\sigma_{i}\rightarrow\sigma_{i}/N_{\pm}$ and define the vortex number density and vortex charge density (vorticity) on a curved surface as following:
$n_{\pm}(x^{\mu})$ $=\frac{1}{N_{\pm}\sqrt{\det g_{\mu\nu}}}\sum^{N_{\pm}}_{i}\delta(x^{\mu}-x^{\mu}_{i}),$ (7)
$\sigma(x^{\mu})$ $=\frac{1}{N_{\pm}\sqrt{\det g_{\mu\nu}}}\sum_{i}\sigma_{i}\delta(x^{\mu}-x^{\mu}_{i}).$ (8)
Here $n_{\pm}$ satisfy the normalization condition $\int{\rm d}\Omega\,n_{\pm}=1$ and ${\sigma}(x^{\mu})=n_{+}(x^{\mu})-n_{-}(x^{\mu})$. In the following we consider vortex distributions at large scales and treat $\sigma$, $n_{\pm}$ and $\psi$ as smooth functions at scales much larger than the mean seperation $\ell$.
For given vortex distributions $n_{\pm}$, the von Neumann entropy reads Joyce and Montgomery (1973)
$S=-\int{\rm d}\Omega\left(n_{+}\log n_{+}+n_{-}\log n_{-}\right),$ (9)
and the energy
$E=\frac{1}{2}\int{\rm d}\Omega\ |\nabla\psi(\textbf{r})|^{2}=\frac{1}{2}\int{\rm d}\Omega\ \psi(\textbf{r}){\sigma}(\textbf{r}),$ (10)
where we have used the relation
$\nabla^{2}\psi(\mathbf{r})$ $=-{\sigma}(\mathbf{r}).$ (11)
The most probable vortex density distribution is given by maximizing the entropy Eq. (9) subject to given energy $E$ and vortex number $N_{\pm}$:
$\delta S-\beta\delta E-\mu_{+}\delta N_{+}/N_{+}-\mu_{-}\delta N_{-}/N_{-}=0,$ (12)
where $\beta,\mu_{\pm}$ are Lagrange multipliers and have the interpretation of inverse temperature and chemical potentials in microcanonical ensemble, respectively. The variation equation (12) gives rise
$n_{\pm}$ $=\exp[\mp\beta\psi(r)+\gamma_{\pm}],$ (13)
[…]
$\nabla^{2}\psi(\mathbf{r})=\exp[\beta\psi(r)+\gamma_{-}]-\exp[-\beta\psi(r)+\gamma_{+}].$ (14)
IV Onsager vortex clusters on a sphere
In this work, we focus on Onsager vortex clusters on a sphere.
IV.1 Stereographic coordinates
The Riemannian metric on a sphere is
$ds^{2}=R^{2}d\theta^{2}+R^{2}\sin^{2}\theta d\phi^{2},$ (15)
where $R$ is the radius, $\theta$ is the polar angle and $\phi$ is the azimuthal angle. In spherical coordinates, the fluid velocity $\textbf{u}=\nabla\psi\times\textbf{e}_{r}$, where $\textbf{e}_{r}$ is the radial unit vector.
The Cartesian coordinates relate to the spherical coordinates via
$\xi=R\sin\theta\cos\phi,\quad\eta=R\sin\theta\sin\phi,\quad\zeta=R\cos\theta.$ (16)
The projection from the north pole to the $\zeta=0$ plane induces stereographic coordinates $z=x^{1}+ix^{2}$, where $x^{1}$ and $x^{2}$ axes are identified as the $\xi$ and $\eta$ axes, respectively (Fig. 1). For a sphere, stereographic coordinates are isothermal coordinates and are related to the spherical coordinates by
$z=R\cot\left(\frac{\theta}{2}\right)e^{i\phi}.$ (17)
In terms of stereographic coordinates, the metric reads
$h=\frac{4R^{4}}{(R^{2}+|z|^{2})^{2}}.$ (18)
It is sometimes convenient to introduce polar coordinates $z=\rho e^{i\phi}$, where $\rho=\cot(\theta/2)$ and the polar angle $\phi$ is chosen to coincide with the azimuthal angle. The relations between the three coordinate systems are
$\xi$ $=$ $\frac{2x^{1}R^{2}}{R^{2}+|z|^{2}}=\sqrt{h}x^{1}=R\sin\theta\cos\phi,$ (19)
$\eta$ $=$ $\frac{2x^{2}R^{2}}{R^{2}+|z|^{2}}=\sqrt{h}x^{2}=R\sin\theta\sin\phi,$ (20)
$\zeta$ $=$ $R\frac{R^{2}-|z|^{2}}{R^{2}+|z|^{2}}=R\cos\theta,$ (21)
and we use Cartesian, spherical or stereographic coordinates wherever convenient.
Figure 1: Stereographic projection from the north pole. The point ${\rm P}(\theta,\phi)$ on the southern hemisphere is projected onto the $\zeta=0$ plane and is denoted as ${\rm P}^{\prime}$.
In stereographic coordinates, the Green’s function reads Kimura (1999)
$G(z,z_{i})$ $=-\frac{1}{2\pi}\log\frac{|z-z_{i}|}{\sqrt{|z-z_{i}|^{2}+|z\bar{z}_{i}+R^{2}|}}$ (22)
and $R_{\rm robin}=(1/2\pi)\log 2$. In spherical coordinates,
$G(\theta,\phi;\theta_{i},\phi_{i})$ $=-\frac{1}{2\pi}\log\sqrt{2R^{2}-2\mathbf{r}\cdot\mathbf{r}_{i}}$ (23)
$=-\frac{1}{2\pi}\log\left(\sqrt{2}R\sqrt{1-\cos\Theta}\right)$
$=-\frac{1}{2\pi}\log\left(2R\sin\frac{\Theta}{2}\right)$
where $\Theta$ is the angle between $\mathbf{r}$ and $\mathbf{r}_{i}$, and $\cos\Theta=\cos\theta\cos\theta_{i}+\sin\theta\sin\theta_{i}\cos(\phi-\phi_{i})$. It is interesting to notice that $\sqrt{2R^{2}-2\mathbf{r}\cdot\mathbf{r}_{i}}$ is the Euclidean distance, but not the geodesic distance, between two points on the sphere.
IV.2 Onset of clustering
The onset of clustering involves global eigen-modes of the Laplace-Bertrami operator. It is then convenient to use spherical coordinates. Hereafter we set $R=1$ for convenience. On a sphere, besides the energy and the number of vortices, the angular momentum $\mathbf{L}=\int\mathbf{r}\times\mathbf{u}$ is also a conserved quantity. In terms of collective variables of vortices, the components of the angular momentum read
$L_{\xi}$ $=\int{\rm d}\Omega\,\sigma\sin\theta\cos\phi,$ (24)
$L_{\eta}$ $=\int{\rm d}\Omega\,\sigma\sin\theta\sin\phi,$ (25)
$L_{\zeta}$ $=\int{\rm d}\Omega\,\sigma\cos\theta.$ (26)
Hereafter we choose the $\zeta$-axis in the direction of total angular momentum and hence $\mathbf{L}=L_{\zeta}$.
Taking into account the conversation of the angular momentum, Eq.(12) becomes
$\delta S-\beta\delta E-\alpha\delta L_{\zeta}-\mu_{+}\delta N_{+}/N_{+}-\mu_{-}\delta N_{-}/N_{-}=0,$ (27)
where $\alpha$ is the Lagrange multiplier associated with the angular momentum and $\omega\equiv\alpha/\beta$ has the meaning of rotation frequency. Accordingly, the vortex number density turns to be
$n_{\pm}=\exp\left[\mp\beta\psi(\textbf{r})\mp\alpha\cos\theta+\gamma_{\pm}\right].$ (28)
Non-uniform solutions at negative temperatures to Eq. (28) describe Onsager vortex states.
Let us consider a solution to Eq. (28) for the given values of $E$ and $L_{\zeta}$, and a nearby solution $n_{\pm}+\delta n_{\pm}$ at $E+\delta E$ and $L_{\zeta}+\delta L_{\zeta}$. The corresponding changes in the constraints are
00 0 $=\int{\rm d}\Omega\ \delta n_{\pm},$ (29)
$\delta E$ $=\int{\rm d}\Omega\ \psi\delta\sigma+\frac{1}{2}\int{\rm d}\Omega\,\delta{\sigma}\delta\psi,$ (30)
$\delta L_{\zeta}$ $=\int{\rm d}\Omega\ \delta\sigma\cos\theta\,.$ (31)
Combining the variations of Eqs. (11) and (28) in leading order, we obtain
$\nabla^{2}\delta\psi$ $=-\delta{\sigma}(\textbf{r})$
$=n\psi\delta\beta+n\beta\delta\psi-n_{+}\delta\gamma_{+}+n_{-}\delta\gamma_{-}+n\cos\theta\delta\alpha,$
where $n=n_{-}+n_{+}$ is the total density of vortices, and
$\nabla^{2}=\frac{1}{\sin\theta}\frac{\partial}{\partial\theta}\left(\sin\theta\frac{\partial}{\partial\theta}\right)+\frac{1}{\sin^{2}\theta}\frac{\partial^{2}}{\partial\phi^{2}}$ (33)
is the Laplace-Bertrami operator in spherical coordinates.
We consider vortex clustered states near the homogeneous state $n_{\pm}=n_{0}=1/4\pi$. For the homogeneous state $\alpha=0$ and $\sigma=\psi=0$, then Eqs. (29),(30) and (31) become
$0=$ $\delta\gamma_{+}+\delta\gamma_{-},$ (34)
$0=$ $\delta\gamma_{+}-\delta\gamma_{-}-2\beta n_{0}\int{\rm d}\Omega\,\delta\psi,$ (35)
$\delta E=$ $\frac{1}{2}\int{\rm d}\Omega\,\delta{\sigma}\,\delta\psi,$ (36)
$\delta L_{\zeta}=$ $-2n_{0}\beta\int{\rm d}\Omega\cos\theta\,\delta\psi-\frac{2}{3}\delta\alpha,$ (37)
where $\delta\alpha$ is assumed to be the same order as $\delta\psi$. Let us introduce operator $\mathcal{P}$:
$\mathcal{P}\delta\psi$ $\equiv\nabla^{2}\delta\psi-2n_{0}\bigg{[}\beta\delta\psi-n_{0}\beta\int d\Omega\,\delta\psi$ (38)
$-3n_{0}\beta\cos\theta\int d\Omega\cos\theta\delta\psi-\frac{3}{2}\cos\theta\delta L_{\zeta}\bigg{]}.$
Then Eq. (IV.2) becomes a zero mode equation of the operator $\mathcal{P}$, i.e.,
$\mathcal{P}\delta\psi=0.$ (39)
In order to solve the zero mode equation, we consider an expansion
$\delta\psi$ $=\sum^{\infty}_{\ell=0}\sum^{\ell}_{m=-\ell}\epsilon f_{\ell m}\psi_{\ell m}(\theta,\phi),$ (40)
where $\psi_{\ell m}$ satisfies $\mathcal{P}\psi_{\ell m}=0$, $f_{\ell m}$ is the mode coefficient, and $\epsilon\ll 1$ is a small amplitude. We denote $\delta L=\epsilon L_{0}$, $\delta\alpha=\epsilon\beta\omega$, and $\delta E=\epsilon^{2}E_{0}$.
We find that
$\psi_{\ell m}(\theta,\phi)=c_{\ell m}{\rm Re}({\rm Y}_{\ell m})+b\cos\theta$ (41)
solves Eq. (39), if
$\beta=-\frac{\ell(\ell+1)}{2n_{0}},$ (42)
and
$b=-\frac{n_{0}\beta\omega}{1+n_{0}\beta}$ (43)
for $\ell\neq 1$. While for $\ell=1$, $b$ can be arbitrarily and we choose that $b=0$ without losing generality. Here ${\rm Y}_{\ell m}$ is the spherical harmonic function. For weekly clustered states, the vorticity field $\sigma=\sigma_{\ell m}\equiv\delta\sigma=\epsilon\ell(\ell+1)\psi_{\ell m}$.
Given the values of $E_{0}$ and $L_{0}$, $c_{\ell m}$ and $\omega$ are determined by
$E_{0}$ $=-\frac{1}{2}\int{\rm d}\Omega\,\psi_{\ell m}(\theta,\phi)\nabla^{2}\psi_{\ell m}(\theta,\phi),$ (44)
$L_{0}$ $=-2n_{0}\beta\int{\rm d}\Omega\cos\theta\ \psi_{\ell m}-\frac{2}{3}\beta\omega.$ (45)
For $\ell=m=0$, $E_{0}=0$ and $L_{0}=0$, hence only modes with $\ell>0$ are relevant. In the following we focus on rotation free states, namely $\omega=0$, and in this case $\psi_{\ell m}=c_{\ell m}\rm{Re}({\rm Y}_{\ell m})$.
IV.2.1 Sandwich state: clusters with zero angular momentum
In this subsection, we consider clustered states with zero angular momentum $L_{0}=0$. Clearly modes with $\ell=2$ have higher statistical weight (larger entropy) than the modes with $\ell>2$.
For $m=0$,
$\psi_{20}=c_{20}{\rm Y}_{20}=\frac{c_{20}}{4}\sqrt{\frac{5}{\pi}}(3\cos^{2}\theta-1),$ (46)
where $c_{20}=\sqrt{E_{0}/3}$. This mode describes the clustering of positive (negative) vortices around the poles and negative (positive) vortices around the equator [Fig.2 (a)]. We refer to this state as sandwich state. The onset of the sandwich state occurs at $\beta=\beta_{c}=-12\pi$.
For $m=1,2$,
$\psi_{22}(\theta,\phi)$ $=c_{22}{\rm Re}({\rm Y}_{22})=\frac{c_{22}}{4}\sqrt{\frac{15}{2\pi}}\cos(2\phi)\sin^{2}\theta,$ (47)
$\psi_{21}(\theta,\phi)$ $=c_{21}{\rm Re}({\rm Y}_{21})=-\frac{c_{21}}{4}\sqrt{\frac{15}{2\pi}}\cos\phi\sin 2\theta,$ (48)
where $c_{22}=c_{21}=\sqrt{2E_{0}/3}$. It happens that these two modes are connected by a $\textrm{SO}(3)$ rotation, i.e., ${\rm Y_{21}}(\mathbf{r}^{\prime})={\rm Y_{22}}(\mathbf{r})$, where $\mathbf{r}^{\prime}=R(-\pi/4,\pi/2,\pi/2)\,\mathbf{r}$ and $R(\theta_{1},\theta_{2},\theta_{3})=R_{\zeta}(\theta_{3})R_{\eta}(\theta_{2})R_{\zeta}(\theta_{1})$ is the $\textrm{SO}(3)$ rotation operator and $\{\theta_{i=1,2,3}\}$ are the Euler angles. Therefore, modes $\psi_{22}$ and $\psi_{21}$ are degenerate and describe the same state. We refer to this state as quadrupole state [Fig.2(b)].
Figure 2: Vorticity $\sigma$ for the sandwich (a), the quadrupole (b) and the dipole (c) state. The extreme values of $\sigma$ are scaled to unity for convenience.
For states with $\ell=2$, the dipole moment defined as
[…]
vanishes. In order to quantify the clustered states with zero angular momentum, it is necessary to introduce the quadrupole tensor
[…]
where $\{\textbf{e}_{i=\xi,\eta,\zeta}\}$ are unit vectors in Cartesian coordinates. The quadrupole moment is defined as
$Q[\sigma]\equiv\sqrt{q^{2}_{1}+q^{2}_{2}+q^{2}_{3}},$ (51)
where $\{q_{1},q_{2},q_{3}\}$ are eigenvalues of $\mathcal{Q}_{ij}$. We denote $Q_{s}$ and $Q_{q}$ as the quadrupole moment for the sandwich and the quadrupole state, respectively.
For the sandwich state, $\ell=2,m=0$,
$\mathcal{Q}[\sigma_{20}]=\epsilon c_{20}\sqrt{\frac{36\pi}{5}}\left(\begin{array}[]{ccc}-1&0&0\\ 0&-1&0\\ 0&0&2\\ \end{array}\right)$ (55)
and its eigenvalues are $\epsilon c_{20}\sqrt{36\pi/5}\,\{-1,-1,2\}$. The quadrupole momentum for the sandwich state is
$Q_{s}=Q[\sigma_{20}]=Q_{0}|E-E_{c}|^{\nu},$ (56)
where $\nu=1/2$, $Q_{0}=\sqrt{216\pi/15}$ and $E_{c}=0$. For quadrupole states, the quadrupole tensors are
$\mathcal{Q}[\sigma_{21}]$ $=-\epsilon{c_{21}}\sqrt{\frac{54\pi}{5}}\left(\begin{array}[]{ccc}0&0&1\\ 0&0&0\\ 1&0&0\\ \end{array}\right){,}$ (60)
and
$\mathcal{Q}[\sigma_{22}]$ $=\epsilon{c_{22}}\sqrt{\frac{54\pi}{5}}\left(\begin{array}[]{ccc}1&0&0\\ 0&-1&0\\ 0&0&0\\ \end{array}\right).$ (64)
Since $\sigma_{21}(R\mathbf{r})=\sigma_{22}(\mathbf{r})$, it is easy to see that $\mathcal{Q}[\sigma_{21}]=R\mathcal{Q}[\sigma_{22}]R^{T}$, where $R=R(-\pi/4,\pi/2,\pi/2)$. Their common eigenvalues are $-{c_{21}}\epsilon\sqrt{54\pi/5}\,\{-1,0,1\}$. Moreover the quadrupole momentum for the quadrupole state takes the same value as for the sandwich state, namely, $Q_{q}=Q[\sigma_{21}]=Q[\sigma_{22}]=Q[\sigma_{20}]=Q_{s}$.
Using Eq. (9), we find that the entropy of the sandwich state is larger than the quadrupole state,
$S[\sigma_{20}]>S[\sigma_{21}]=S[\sigma_{22}].$ (65)
Hence under condition of zero angular momentum $L_{0}=0$, the sandwich state has the highest statistical weight. Although this conclusion is made based on the perturbative analysis at low energies, it will be confirmed by later analysis at high energies and Monte Carlo simulations at moderate energies.
IV.2.2 dipole states: vortex clustered states with finite angular momentum
We now consider states with finite angular momentum $L_{0}\neq 0$. For $\ell=1,m=0$, the corresponding stream function is
$\psi_{10}=c_{10}{\rm Y}_{10}=c_{10}\sqrt{\frac{3}{4\pi}}\cos\theta,$ (66)
where $c_{10}=\sqrt{E_{0}}$.
The dipole moment for this state does not vanish and its magnitude reads
$D=|\mathbf{D}|=4\sqrt{\frac{\pi}{3}}\sqrt{E}.$ (67)
We refer this state as dipole state [Fig. 2 (c)] and the onset of the dipole state occurs at $\beta=\beta^{d}_{c}=-4\pi$. For the dipole state, the angular momentum is $L_{\zeta}=4\sqrt{\pi/3}\sqrt{E}$. Note that the vorticity distribution for the dipole state at low energies is identical to the vorticity distribution for a stationary vortex flow on a sphere Xiong and Yu (2024).
IV.3 High energy configurations
In this subsection, we consider the sandwich state and the dipole state at high energies. In the high energy limit, vortices and antivortices are well separated and are concentrated in small regions of a sphere. It is then convenient to use the stereographic projection method to find approximate density distributions of vortices on a sphere.
In stereographic coordinates Eq. (14) becomes
$\ (1+z\bar{z})^{2}\partial_{z}\partial_{\bar{z}}\psi=e^{\beta\psi(z,\bar{z})+\gamma_{-}}-e^{-\beta\psi(z,\bar{z})+\gamma_{+}}$ (68)
for rotation free states ($\omega=0$). The stereographic projection maps the southern hemisphere onto a disc on the $\zeta=0$ plane (Fig.1) and transforms vortices in the region around southern pole and vortices in the region around the equator to the region around the origin and the region near the boundary of the disc, respectively.
IV.3.1 Sandwich states at high energies
In the high energy limit, the interaction between vortices near the poles and antivortices near the equator can be neglected. Let us firstly consider the region $|z|\ll 1$, in which Eq. (68) can be well approximated as
$\partial_{z}\partial_{\bar{z}}\psi=-e^{-\beta\psi(z,\bar{z})+\gamma_{+}}.$ (69)
For axisymmetric distributions, Eq. (69) becomes
$\frac{1}{4}\frac{1}{\rho}\frac{{\rm d}}{{\rm d}\rho}\left(\rho\frac{{\rm d}}{{\rm d}\rho}\psi\right)=-e^{-\beta\psi(\rho)+\gamma_{+}},$ (70)
and has the following exact solution
$\psi(\rho)$ $=-\frac{2}{\beta}\log\left(\frac{2}{\beta\rho^{2}e^{\gamma_{+}}-2}\right)$ (71)
with boundary condition $\psi(\rho=0)=\psi^{\prime}(\rho)|_{\rho=0}=0$ Xiong et al. (2023). The vortex density reads
$n_{+}(\rho)=\exp(-\beta\psi+\gamma_{+})=\frac{4e^{\gamma_{+}}}{(\beta\rho^{2}e^{\gamma_{+}}-2)^{2}}.$ (72)
For the sandwich state, since only half positive vortices are distributed around the south pole (the other half positive vortices are distributed around the north pole), the normalization condition is
[…]
which gives rise the relation between $e^{\gamma_{+}}$ and $\beta$ [Fig. 10(a) in Appendix]. When $\beta\rightarrow\beta_{s}=-16\pi$, vortices around the pole collapse into a point, known as the supercondensation. Hence the parameter range for the sandwich state is $\beta_{s}<\beta<\beta_{c}$. The limit configuration at $\beta=\beta_{s}$ is $n_{+}(\theta)=[\delta(\theta)+\delta(\theta-\pi)]/(4\pi\sin\theta)$. It is worthwhile mentioning that the value of $\beta_{s}$ depends on the normalization condition. For vortices confined to a disc on a plane, the vortex density profile is the same as Eq. (72), however the vortex number is normalized to unity, which gives rise that $\beta_{s}=-8\pi$ (in convention of this paper) Smith and O’Neil (1990); Yu et al. (2016); Kiessling (1995); Patil and Dunkel (2021); Xiong et al. (2023).
In spherical coordinates, the stream function and the vortices density on the southern hemispheres [$\theta\in(\pi/2,\pi)$] are [see also Fig. 3(a)]
$\psi(\theta)$ $=-\frac{2}{\beta}\log\left[\frac{2}{\beta\cot^{2}\left(\frac{\theta}{2}\right)e^{\gamma_{+}}-2}\right],$ (74)
$n_{+}(\theta)$ $=\frac{4e^{\gamma_{+}}}{\left[\beta\cot^{2}\left(\frac{\theta}{2}\right)e^{\gamma_{+}}-2\right]^{2}}.$ (75)
Figure 3: The vortex density distributions for the sandwich state at high energies. (a) The distribution of positive vortices near poles [Eq. (75)] . (b) The distribution of negative vortices near the equator at $\beta=\beta_{s}$ [Eq. (82)].
Now we consider the distribution of negative vortices in the region near the equator which is mapped to the region near the edge of the disc. For $|z|\sim 1$, Eq. (68) can be well approximated as
$4\partial_{z}\partial_{\bar{z}}\psi$ $=e^{\beta\psi(z,\bar{z})+\gamma_{-}}.$ (76)
For axis-symmetric solutions, Eq. (76) becomes
$\frac{1}{\rho}\frac{{\rm d}}{{\rm d}\rho}\left(\rho\frac{{\rm d}}{{\rm d}\rho}\psi\right)$ $=e^{\beta\psi(\rho)+\gamma_{-}},$ (77)
and admits an exact solution
$\psi(\rho)=\frac{2}{\beta}\log\left(\frac{2A\rho^{(-1+A/2)}}{2+A+(A-2)\rho^{A}}\right),$ (78)
with the boundary condition Xiong et al. (2023)
$\psi(\rho=1)=\psi^{\prime}(\rho)|_{\rho=1}=0.$ (79)
Here $A=\pm\sqrt{4-2\beta e^{\gamma_{-}}}$. As it follows, the density is
$n_{-}(\rho)=\frac{4A^{2}\rho^{(A-2)}e^{\gamma_{-}}}{\left[2+A+(A-2)\rho^{A}\right]^{2}}.$ (80)
The relation between $e^{\gamma_{-}}$ and $\beta$ is given by the normalization condition Eq. (73) ($n_{+}\rightarrow n_{-}$). Moreover, the smooth condition $n^{\prime}_{-}(\rho)|_{\rho=0}=0$ requires $e^{\gamma_{-}}\beta<-5/2$ which gives rise to that $\beta<-7.11\pi$ [Fig. 10 (b) in Appendix]. Hence for $\beta$ within the range $\beta_{s}<\beta<\beta_{c}$, the smooth condition is automatically satisfied.
In spherical coordinates, the stream function and the vortex density distribution of negative vortices near the equator [$\theta\in(\pi/2,\pi)$] read
$\psi(\theta)$ $=\frac{2}{\beta}\log\left[\frac{2A\cot(\theta/2)^{(-1+A/2)}}{2+A+(A-2)\cot(\theta/2)^{A}}\right],$ (81)
$n_{-}(\theta)$ $=\frac{4A^{2}\cot^{(A-2)}(\theta/2)e^{\gamma_{-}}}{\left[2+A+(A-2)\cot^{A}(\theta/2)\right]^{2}}.$ (82)
Induced by the boundary condition Eq. (79), the fluid velocity $u_{\phi}=-\partial\psi/\partial\theta$ on the equator vanishes, namely $u_{\phi}(\theta=\pi/2)=\psi^{\prime}(\rho)\rho^{\prime}(\theta)|_{\theta=\pi/2}=0$, which is consistent with the fluid velocity field on the equator at low energies [$\psi_{20}^{\prime}(\theta)|_{\theta=\pi/2}=0$]. At $\beta=\beta_{s}$, $e^{\gamma_{-}}\approx 0.167,A\approx 4.556$ [Fig. 10(b) in Appendix], the vorticity for the sandwich state reads $\sigma_{s}(\textbf{r})=\Big{[}\delta(\theta)+\delta(\theta-\pi)\Big{]}/(4\pi\sin\theta)-n_{-}(\theta),$ and the quadrupole moment reaches the maximum value $Q^{\rm max}_{s}\approx 1.59$. It is easy to see that the quadrupole state also reaches the high energy limit at $\beta=\beta_{s}$, and the corresponding vorticity is ${\sigma}_{q}(\textbf{r})=\Big{[}\delta(\theta-\pi/4,\phi)+\delta(\theta-3\pi/4,\phi-\pi)-\delta(\theta-3\pi/4,\phi)-\delta(\theta-\pi/4,\phi-\pi)\Big{]}/(2\sin\theta)$ which gives rise the maximum quadrupole moment $Q^{\rm max}_{q}=3\sqrt{2}/2>Q^{\rm max}_{s}$. Note that for the quadrupole state locations of the clusters at supercondensation coincide with mechanical equilibrium positions for $4$ point vortices on a sphere.
IV.3.2 Dipole states at high energies
At high energies, for the dipole state, vortices with opposite sign are concentrated near the poles. The density distribution of positive vortices near the southern pole is identical to Eq. (72) and the only difference is that here the vortex density is normalized to unity but not 1/2. Hence for the dipole state, the supercondensation occurs at $\beta=\beta^{d}_{s}=-8\pi$ which is the same as what for vortices confined to a disc on a plane Smith and O’Neil (1990); Yu et al. (2016); Xiong et al. (2023). In spherical coordinates, vortex densities are [see also Fig. 4]
$n_{-}(\theta)$ $=\frac{4e^{\gamma_{-}}}{\Big{[}\beta\cot^{2}(\theta/2)e^{\gamma_{-}}-2\Big{]}^{2}}\quad\theta\in(\pi/2,\pi),$ (83)
$n_{+}(\theta)$ $=\frac{4e^{\gamma_{+}}}{\Big{[}\beta\tan^{2}(\theta/2)e^{\gamma_{+}}-2\Big{]}^{2}}\quad\theta\in(0,\pi/2),$ (84)
and the corresponding angular momentum is
$L_{\zeta}(\beta)=$ $\frac{8\pi e^{\gamma_{-}}}{(2+\beta e^{\gamma_{-}})^{4}}\Big{[}\beta e^{\gamma_{-}}\Big{(}\beta e^{\gamma_{-}}(-2\beta+5+32\log 2)$ (85)
$+20-32\log 2\Big{)}-16\beta e^{\gamma_{-}}(\beta e^{\gamma_{-}}-1)\log(2-\beta e^{\gamma_{-}})+4\Big{]}.$
Figure 4: The total vortex density for the dipole state at high energies [Eqs. (83) and (84)].
When $\beta\rightarrow\beta^{d}_{s}$, $n_{-}=\delta(\theta-\pi)/(2\pi\sin\theta),n_{+}=\delta(\theta)/(2\pi\sin\theta)$, and the angular momentum reaches the upper bound $L_{\zeta}=L^{\rm{max}}_{\zeta}=2$.
For the dipole state at low energies, the fluid velocity $u_{\theta}=0$, $u_{\phi}\propto\sin\theta$ and $u_{\phi}$ vanishes at poles (Fig. 5). At high energies, the velocity field is
$u_{\phi}(\theta)=-\frac{\partial\psi}{\partial\theta}=-\frac{2\cot(\theta/2)\gamma_{-}}{[-2+\beta\gamma_{-}\cot^{2}(\theta/2)]\sin^{2}(\theta/2)},$ (86)
where we have used Eq. (74). Consistently, it also vanishes at poles. At low energies, fluid velocity field $u_{\phi}$ reaches the peak value on the equator ($\theta=\pi/2$). While at high energies, two peaks of the velocity field start to develop and move towards to poles as increasing the energy (Fig. 5). This is because for tightly clustered vortices around the poles, the fluid velocity is considerably large on the edges of the vortex clusters which locate between the poles and the equator.
Figure 5: The velocity field $u_{\phi}$ for the dipole state at a low energy (orange solid line) and at a high energy (blue solid line). The parameter region for the dipole state is $-8\pi=\beta^{d}_{s}<\beta<\beta^{d}_{c}=-4\pi$. When $\beta\to\beta^{d}_{s}$, $u_{\phi}(\theta=\pi/2)\rightarrow 1/2\pi$. Note that here the fluid velocity field on the northern hemisphere is obtained by shifting the coordinate of Eq. (86) ($\theta\rightarrow\pi-\theta$). The two expressions work well only near the poles and take different values on the equator. However the difference decreases as increasing energy and for the $\beta$ value we choose here, the discrepancy is invisible.
The quadrupole state at high energies can be easily constructed from Eqs. (83) and  (84). From the approximate high energy expressions of vortex densities, we can calculate the entropy and find again that $S[\text{sandwich state}]>S[\text{quadrupole state}]$.
IV.3.3 General solutions
A general solution to Eq. (69) is available Crowdy (1997); Henrici (1993) :
$\psi(z,\bar{z})=$ $-\frac{1}{\beta}\log\left[\frac{2f^{\prime}(z)\bar{f}^{\prime}(\bar{z})}{-c\beta(1+f(z)\bar{f}(\bar{z}))^{2}}\right],$ (87)
$n(z,\bar{z})=$ $c\exp(-\beta\psi),$ (88)
where $f(z)$ is a meromorphic function of $z$ in a simply connected domain and $\bar{f}(\bar{z})$ is the complex conjugate of $f(z)$. Furthermore , it is required that $f(z)$ has at most simple poles and $f^{\prime}(z)\neq 0$ in the consider domain Henrici (1993); CROWDY (2004).
Here we show that Eqs. (71) and (78) fall into this general solution by specifying function $f(z)$ properly. Let us firstly consider the vortex distribution around the pole. For an axisymmetric distribution, the stream function $\psi(z,\bar{z})$ must be the function of $z\bar{z}$ or $\rho$ only. Hence $f(z)$ has a unique form and $f(z)=az^{b}$, here $a\in\mathbb{C},b\in\mathbb{Z}$. The boundary condition $\psi(\rho=0)=\psi^{\prime}(\rho)|_{\rho=0}=0$ determines that $a=\sqrt{-c\beta/2}$, $b=1$, and $c=e^{\gamma_{+}}$, which leads to  Eq. (71) precisely.
For the distribution near the edge (or equator in spherical coordinates), the general solution Eq. (87) gives rise to Eq. (78) by choosing $f(z)=\ \sqrt{(A-2)/(A+2)}\ z^{A/2}$, where $A=\pm\sqrt{4-2\beta e^{\gamma_{-}}}$, and $c=e^{\gamma_{-}}/4$. Here the boundary condition $\psi(\rho=1)=\psi^{\prime}(\rho)|_{\rho=1}=0$ is fulfilled. However it is important to notice that the exponent $A/2$ is not an integer in general. Such a choice of $f(z)$ may not be a meromorphic function and it does not meet the requirements for Eq. (87). As shown previously, such a solution is indeed the solution to Eq. (77), suggesting that the validity condition of Eq. (88) can be extended.
IV.4 Gaussian clustered states
In this section, we discuss a limit distribution when $\beta\rightarrow 0$, $\omega\rightarrow\infty$ simultaneously while $\alpha=\beta\omega$ is finite. In this limit, Eq. (28) admits an exact solution:
$n_{\pm}(\theta)=$ $\frac{\alpha}{4\pi\sinh\alpha}\exp(\mp\alpha\cos\theta),$ (89)
$\psi(\theta)=$ $-\frac{1}{8\pi\sinh\alpha}\Big{[}e^{\alpha}{\rm Ei}\Big{(}\alpha(\cos\theta-1)\Big{)}-e^{\alpha}{\rm Ei}\Big{(}-\alpha(\cos\theta+1)\Big{)}$
$+e^{-\alpha}{\rm Ei}\Big{(}\alpha(1-\cos\theta)\Big{)}-e^{-\alpha}{\rm Ei}\Big{(}\alpha(\cos\theta+1)\Big{)}$
$-2\alpha C\ \rm{arctanh}(\cos\theta)\Big{]},$ (90)
where $\alpha\in(-\infty,\infty)$, ${\rm Ei}(x)=\int^{x}_{-\infty}e^{t}/t\ {\rm d}t$, and $C$ is a constant determined by proper smooth conditions. The angular momentum for this Gaussian state is
$L_{\zeta}=\int(n_{+}-n_{-})\cos\theta{\rm d}\Omega=\frac{2}{\alpha}(1-\alpha\coth{\alpha}).$ (91)
The fluid velocity field for this state is
$u_{\theta}$ $=\frac{1}{\sin\theta}\frac{\partial\psi}{\partial\phi}=0,$ (92)
$u_{\phi}$ $=-\frac{\partial\psi}{\partial\theta}=\frac{\alpha}{4\pi\sinh\alpha\sin\theta}\left[\frac{2\cosh({\alpha\cos\theta})}{\alpha}+C\right].$ (93)
Here we impose the zero velocity condition at poles, namely $u_{\phi}(\theta=0)=u_{\phi}(\theta=\pi)=0$, giving rise that $C=-2\cosh(\alpha)/\alpha$. Figure 6 shows the vorticity distribution, the energy and the angular momentum as functions of $\alpha$. Note that the condition $u_{\phi}(\theta=\pi/2)=0$ leads to unphysical results $u_{\phi}(\theta=0)=u_{\phi}(\theta=\pi)=\infty$ for $\alpha\neq 0$.
Figure 6: (a) Vorticity $\sigma=n_{+}-n_{-}$ for the Gaussian state at $\alpha=-8$. (b) and (c) show the energy and the angular momentum of as functions of $\alpha$, respectively.
Similar to the fluid velocity field for the dipole state, for the Gaussian state the peak fluid velocity appears on the equator for small values of $|\alpha|$ and moves towards to poles as $|\alpha|$ exceeds a certain value of $|\alpha_{c}|$ (Fig. 7), where $|\alpha_{c}|$ satisfies ${\alpha_{c}}^{2}-\cosh(\alpha_{c})+1=0$. The amplitude of the fluid velocity on the equator is $|u_{\phi}(\theta=\pi/2)|=(1-1/\cosh\alpha)/(2\pi)$ and reaches the maximum value $1/2\pi$ when $|\alpha|\to\infty$.
Figure 7: The amplitude of the fluid velocity field $u_{\phi}$ for different values of $|\alpha|$.
V Microcanonical Monte Carlo Simulations
In this section, we show numerical results of Monte Carlo (MC) simulations of a large number of point vortices on a unit sphere. The simulations are performed for a range of increasing energies and at fixed angular momentum $L_{\zeta}=0$.
V.1 Sandwich state and the quadrupole moment
In our microcanonical MC simulations, we adopt the general scheme developed in Refs. Smith (1989); Smith and O’Neil (1990); Yu et al. (2016) which are based on the demon algorithm Creutz (1983). In particular, we closely follow the algorithm developed in Ref. Yu et al. (2016) for a neutral vortex point vortex system. We find that the sandwich state emerges as increasing energy. A comparison between the analytical predictions of the mean field theory and the microcanonical MC sampling for point vortices is shown in Fig. 8. The numerical results for the values of $\nu$, $Q_{0}$ and the upper bound of the quadrupole moment are in good agreement with the mean field predictions.
Figure 8: A comparison between the mean field theory and MC simulations for $N=1000$ point vortices. (a) and (b) show results from the mean field theory and Monte Carlo sampling for the scaled vorticity $\tilde{\sigma}$, respectively. Here $\widetilde{\sigma}=\sigma$ for $E=0.005$, $E=0.01$, and $\widetilde{\sigma}=\sigma/20$ for $E=0.1$. The purpose for introducing the scaled vorticity $\tilde{\sigma}$ is to present the large energy range on a single color map. (c) The quadrupole moment $Q$ for increasing energy $E$ (circle red) and the mean field prediction Eq. (56) (solid blue). The quadrupole moment is well bounded by the maximum possible value $Q^{\rm max}_{s}$ predicted by the mean field theory. The inset in (c) shows the $N$-dependence behavior of $Q$ in the vicinity of the clustering transition. For finite $N$, the transition energy $E_{c}$ is finite and approaches to zero as $N\rightarrow\infty$. The finite values of $Q\sim 1/\sqrt{N}$ below the transition energy is nearly energy independent and is due to uncorrelated fluctuations. The data presented here are taken in runs of $40000$ MC steps at each value of energy.
V.2 Sandwich state vs. quadrupole state
Our theory predicts that the sandwich state has higher entropy than the quadrupole state (see Sec. IV.2), for the fixed values of energy, number of vortices, and zero angular momentum. In order to verify this analytical result numerically, we perform the MC sampling starting from a quadrupole configuration at a moderate energy. We find that the vortex distribution saturates at the sandwich state [Fig.9(a)(b)], indicating that indeed the sandwich state has greater statistical weight than the quadrupole state. This can be further verified by measuring directly the entropy of the initial and final distributions for a certain range of energies [Fig.9(c)]. Here the initial quadrupole distribution is obtained through sampling of the quadrupole mode $\sigma_{21}$ (or $\sigma_{22}$).
Figure 9: (a) A point vortex system containing $N=1000$ vortices which is initially prepared at a quadrupole state with $E=0.06$ and $L_{\zeta}=0$. (b) The finial state after $160000$ MC steps and the final state is the sandwich state. (c) A comparison between the entropy of the initial quadrupole state and the final sandwich state. The sphere is divided into $40\times 40$ cells by longitude and latitude (these cells have different area). The entropy is calculated through Eq. (94). Here the entropy is negative because of the artificial choice of the minimum area in the phase space, and only the entropy difference between the two states has physical meaning.
The entropy of a point vortex system confined on a domain can be measured numerically. We divide the domain into a large number of cells with size $a_{i}$, where $i$ denotes the cell index. Each cell contains a sufficient number of vortices. We denote the number of point vortices in a cell of each species as $N_{i,\hat{\sigma}}$ with $\hat{\sigma}=\pm$. The number of states is then given by
$\mathcal{W}\left(\{N_{i,\hat{\sigma}}\}\right)=\prod_{\hat{\sigma}}\left\{\prod_{i}\frac{N_{\hat{\sigma}}!}{N_{i,\hat{\sigma}}!}\left(\frac{a_{i}}{\Omega}\right)^{N_{i,\hat{\sigma}}}\right\},$ (94)
where is $\Omega$ is the area of domain Joyce and Montgomery (1973); Maestrini and Salman (2019). As it follows the entropy per vortex is $S=(1/N)\log\mathcal{W}$. We apply this method to vortices on a unit sphere and compare the entropy between the sandwich state and the quadrupole state [Fig.9(c)].
VI Conclusion and discussion
We investigate Onsager vortex clustered states of point vortices confined on a sphere. We find that for vortex distributions with zero angular momentum, the sandwich state is the most probable state and is characterized by the quadrupole moment tensor. The mean field analytical predictions agree well with the Monte Carlo simulations. It is worthwhile mentioning that Ref. Kanai and Guo (2021) shows that the Gross-Pitaevskii dynamics drives an initial uniform vortex state to the end quadruple state through the evaporative heating process. This result is not necessary inconsistent with the our findings. Since the quadruple distribution also solves the self-consistent equation and hence it is a local maximum entropy state. In the situation described in Ref. Kanai and Guo (2021), vortices may be dynamically trapped in a local equilibrium state. To explore the entropy barrier between the sandwich state and the quadruple state and its effect on the dynamical transitions between the two states deserves future investigations.
Acknowledgement
We thank Y. Xiong and C. Ma for useful discussions. X.Y. acknowledges support from the National Natural Science Foundation of China (Grant No. 12175215), the National Key Research and Development Program of China (Grant No. 2022YFA 1405300) and NSAF (Grant No. U2330401).
Appendix A Appendix
Appendix B Normolization dependence of the onset of the supercondensation
The value of $\beta$ at which the supercondensation occurs depends on the normalization condition. Here we show the method we used to determine the value of $\beta_{s}$ for the sandwich state [see Fig. 10(a)].
Figure 10: (a) The $\beta-e^{\gamma_{+}}$ curve is determined by normalization Eq. (73) and touches the $\beta=\beta_{s}=-16\pi$ curve when the supercondensation occurs. (b) The $\beta-e^{\gamma_{-}}$ curve is determined by the normalization Eq. (73) ($n_{+}\rightarrow n_{-}$). The condition $\beta e^{\gamma_{-}}<-5/2$ ensures $n^{\prime}_{-}(\rho)|_{\rho=0}=0$.